OpenStudy (thatonegirl_):

if f is the antiderivative of the x^2/(1+x^5) such that f(1)=0, then f(4)= ?

OpenStudy (thatonegirl_):

\[\frac{ x^2 }{ 1+x^5 }\]

OpenStudy (helper101):

easy

OpenStudy (thatonegirl_):

help plss i've done it 3 times and can't get the right answer

OpenStudy (helper101):

So you want to compute \int x^2/(1+x^5) dx from 1 to 4. g(x) = x^2 / (1+x^5) is strictly decreasing from 1 to 4 as you can check by dg/dx. And g(1) = 1/2, g(2) = 4/33, g(3) = 9/244, g(4) = 16/1025. If you're allowed to use the calculator, g(1.5) = 0.2618 as well. So (a), (b) are obviously out. (c) is out because the area from 1 to 1.5 is already > 0.2618 * 0.5 = 0.13. Finally, I believe (e) is an overestimate (you'll have to verify it yourself).

OpenStudy (helper101):

I think

OpenStudy (helper101):

does this help

OpenStudy (thatonegirl_):

wait where did u get that answer from?

OpenStudy (helper101):

school teacher y

OpenStudy (thatonegirl_):

lol it was a really fast response and u knew all the answer choices xD

OpenStudy (helper101):

ik right

OpenStudy (thatonegirl_):

so g(x) is what, the original function??

OpenStudy (helper101):

it is in what i said

OpenStudy (helper101):

is this a multiply chose

OpenStudy (helper101):

question

OpenStudy (thatonegirl_):

lol ur answer is straight off yahoo answers xD

OpenStudy (thatonegirl_):

do u know how to do these tho?

OpenStudy (helper101):

yes

OpenStudy (helper101):

i just ask my teacher and he said that

OpenStudy (helper101):

NOT ME

OpenStudy (thatonegirl_):

ok thx for the help

OpenStudy (helper101):

sure np

OpenStudy (helper101):

lol :)